Grade 11General Physics

  1. A projectile is thrown horizontally from the top of a tower and strikes the ground after 3 seconds at an angle of 45 with the horizontal find the height of the tower

Aventika Khemani
8 Years agoGrade 11

1 Answer

Aashu Mittal
8 Years ago
Since initial vertical velocity of horizontal projectile is 0. Using Newton second equation of motion, S=ut+1/2at^2Let H be the height of tower. Therefore, H=1/2gt^2 Now, it is given that t=3 sec. and putting g=10 we will get, H=90/2=45m.........AnsIn this problem, the is not any use of the given angle 45°You can ask me for any query related to this problem.